# Shahriar Noroozizadeh

### v2

@@ B.2.2 Tiny model architectures
For toy experiments on small-scale graphs (§C.3, §D.1.2), including the Tiny Path–Star, Tiny Grid, Tiny Cycle, and Tiny Irregular graphs, we used reduced-size architectures.
Models. We evaluated three model types: (1) a Transformer (TinyGPT), (2) a feed-forward neural network with the same configuration but without residual connection or attention and (TinyNN) (3) a Mamba model (TinySSM) of comparable scale to (1). Exact configurations are described below.
@@
  • For all graphs, all models used a single layer ($N_{\tt layer}=1$), but the model specifications differed across graph types.
@@
  • For Tiny Path–Star and Tiny Grid, TinyGPT used embedding dimension $m_{\tt width}=32$ with $m_{\tt head}=8$, and TinySSM used embedding dimension $8$ with $d_{\text{state}}=2$, $d_{\text{conv}}=2$, and $\text{expand}=1$.
  • For Tiny Cycle, TinyGPT used embedding dimension $16$ with $8$ heads, while TinySSM matched the same specifications as above ($8$-dim, $d_{\text{state}}=2$, $d_{\text{conv}}=2$, $\text{expand}=1$).
  • For Tiny Irregular, TinyGPT again used embedding dimension $16$ with $8$ heads, and TinySSM used embedding dimension $8$ with $d_{\text{state}}=2$, $d_{\text{conv}}=4$, and $\text{expand}=1$.
For TinyNN, we used a large embedding dimension $512$ with $1$ head for the sake of empirical proofs of Section˜3. We elaborate why. First, TinyNN architectures are a clean architecture where one can be certain that freezing the (un)embedding layers results in a purely associative storage; thus they are appropriate for our empirical proofs. Next, in our proofs, we wanted to make two demonstrations. First, that the model memorize geometrically even without bottleneck pressures (as is the case with a large embedding dimension); second, that in sufficiently wide-neck settings, the associative storage is realized within just $2$ learning steps (which we find is true here).
We clarify that even when reducing the embedding dimension of TinyNN to match that of TinyGPT (Tiny Path–Star and Tiny Grid: $m_{\tt width}=32$; Tiny Cycle and Tiny Irregular: $m_{\tt width}=16$), we found that the geometric structure still persisted and that the associative memory variant continued to fully memorize the edge bigrams, albeit with longer optimization steps.
@@ B.3 Training and optimization
Objective. We use next-token cross-entropy over the causal prefix. For first-token-only experiments, the loss is restricted to the first target position. Although we train all our models to $50,000$ epochs, we only need about a couple of thousand epochs to see accuracy gains (e.g., see the per-token accuracy plots in Fig.˜4).
@@
B.4 Evaluation protocols and metrics
@@
  • Per-token accuracy over epochs / Token accuracy: accuracy at each target position, tracked through training (used in Fig. 35).
@@
B.5 Implementation and compute
Code is implemented in PyTorch with standard Transformer components. All runs fit on a single modern GPU (e.g., A100-40GB or A100-80GB for the largest batch-size); per-figure training time depends on $(d,\ell)$ and batch size.
@@ C.1 Path-star task on Mamba SSM
This section demonstrates that the implicit reasoning on path-star graphs observed for Transformers in §2.1, generalizes to the Mamba SSM architecture [60] too. Fig.˜14 is the counterpart of the Transformer accuracy plots in Fig.˜4; Fig.˜15 is the counterpart to Transformer heatmaps in Fig.˜25; Fig.˜16 is the counterpart to the UMAP topology of Transformers in Fig.˜27. A notable difference here is that the Mamba model presents a strong geometry even when trained only on the edges (as seen in Fig.˜15(c)).
Figure 14: (left) Success of in-weights path-star task for Mamba. This figure is a counterpart to Fig.˜4. A next-token-trained Mamba achieves perfect or highly non-trivial accuracy on large path-star graphs $\mathcal{G}_{d,\ell}$ (Section˜A.1). (middle) Learning order of tokens. The tokens of a path are not learned in the reverse order i.e., the model does not learn the right-to-left solution. Thus, gradients from the future tokens are not critical for success (Section˜A.1.1). (right) Success of hardest-token-only task. In fact, the hardest token (the first token) given the leaf is learned in isolation to non-trivial accuracy (Section˜2.2). Success of this $\ell$-fold composition task is hard to explain within the associative memory (§2.3).
@@
(b) Edge-Memorized and
Hardest-Token-Trained
(c) Only Edge-Memorized
Figure 15: Evidence of global geometry in path-star task for Mamba. This figure is a counterpart to Figs.˜6 and 25 for the Mamba SSM architecture. Recall that entry $(i,j)$ is the mean cosine distance between the leaf token in (an unseen) path $i$ (row) and first/hardest token on (an unseen) path $j$ (col). Each heatmap corresponds to a different training objective: Left: trained on edges and path-finding task $(\mathcal{D}_{\texttt{edge}}\cup\mathcal{D}_{\texttt{path}}^{\to})$; Middle: trained on edges and hardest-token-finding task (not presented in the main paper); Right: edges only $(\mathcal{D}_{\texttt{edge}})$. We find that even on these unseen paths, the leaf and first token embeddings cluster together, regardless of whether path-finding supervision exists. Interestingly, compared to the Transformer in Fig.˜25, the Mamba trained only on local supervision (right) exhibits a much stronger geometry.
Figure 16: UMAP projection of token embeddings of Mamba exhibits path-star topology. We corroborate the UMAP [109] observations from the Transformer (Fig.˜27) in Mamba SSM here. Each point is a node embedding; color indicates path identity. Different paths form separated clusters (although this clustering is weaker than for a Transformer); the central node ${v}_{\tt{root}}$ is excluded since it is shared across all paths. The graph has degree $10^{4}$ and path length 66 (including the root). Axes are UMAP components (arbitrary units). Note that we have re-used each color for multiple paths.
@@ C.2 Other large, harder path-finding graphs
Next, we consider variants of the path-finding graph. While the path-star graph is adversarially constructed in certain ways, there is only one decision-making step, which makes planning simpler in a certain way. To make the task harder, we introduce a branching at every node along each path. (Similar tree variants have been considered in [47] for the in-context task, but we are interested in in-weights tasks.)
In the tree-star graph, $\mathcal{T}_{d,\ell}$, there is a central node with degree $d$ and each child node except the leaf has a fixed degree of $2$, as visualized in Fig.˜17(a). The path-length from the root to any leaf is $\ell$. In this graph, two types of learning tasks can be considered, depending on how we split the test and training paths. For a no-overlap setting, we could sample all training paths from one subset of trees, and the test paths from the remaining trees; we call this the split at first token setting, since the test/train split is determined by the first token. A second setting—with some test-train overlap—is one where we reserve some leaves as training goals, and the rest as test-goals. Here, on any test path, a prefix may have participated in a training path. We call this the split at leaf setting. Note that in both variants, all nodes will be sampled as part of the edge-memorization task.
In Fig.˜17(b), we find that on both variants the Transformer achieves non-trivial path-finding accuracy, generalizing our results beyond the path-star task. However, we note that the test-train split at the first token is much harder to succeed at, likely due to no overlaps.
@@ C.3 Tiny graphs
Besides the large path-star graph (of §2) and tree-star graphs (of C.2), we also report the embeddings learned on four tinier graphs for various architectures (details in B.2.2); the figures here are an extension of the Transformer embeddings in Fig.˜1. In these experiments, we train the models purely on local supervision (the edges, presented in both directions) to $100\%$ edge-memorization accuracy: for each vertex, we ensure that its $d$ neighbors appear in the top $d$ softmax probabilities. The graphs include (a) a tiny path-star graph with four paths of length $4$, (b) a $4\times 4$ grid graph, (c) a $15$-node cycle graph and (d) an irregular graph with two asymmetric components. Each visualization includes the embeddings from: an associative memory model (implemented with a neural network with only one trainable matrix sandwiched between (un)embedding layers), a Node2Vec model, the eigenvectors of the graph Laplacian, a Transformer’s token embeddings, a neural network’s first layer, and a Mamba SSM’s token embeddings. For the Node2Vec model we use the top eigenvectors, and for the rest we use UMAP [109] to choose the top directions.
We note that these visualizations by themselves do not guarantee the presence of global information (a model that embeds each node as its adjacency row can still show similar geometries). As a rigorous test, we provide node-node cosine-similarity heatmaps to examine whether it exhibits multi-hop information. These are in Figs.˜24 and 23.
We consolidate the various observations from these figures (Figs.˜18, 19, 20, 21, 24, 23 and 22) and our memorization vs. time plots in Fig.˜7 here:
In the tiny graphs of Figs.˜18, 19, 20 and 21, on various architectures (Node2Vec, graph spectrum, Transformer, neural network, Mamba SSM), we find that:
@@
  5. 5.
    Associative storage can be quickly found by gradient descent (Fig.˜7), as claimed in Section˜3.2, whereas geometric memorization takes longer to discover (Fig.˜22)
@@
    We note that similar geometries arise even with only one direction presented; see §D.1.2. This is the scenario where both types of storage have the same bit and $\ell_{2}$ norm complexity; thus, there are no straightforward implicit pressures that encourage the geometry (Section˜3.2).
@@
Figure 21: Tiny irregular graph: Geometries of various architectures on a small irregular graph of two connected components, both asymmetric. See Section˜C.3. Note that unlike in the other graphs, we do not use a 11-layer model here, but a 33-layered one.
@@
Figure 22: For various learning rates, geometric memorization takes much longer for gradient descent to discover: This is an extended version of Fig.˜8 for two more learning rates. Recall that for the same TinyNN architecture as in Fig.˜7 but without frozen embeddings, we roughly quantify geometric memorization over time via cosine distance across all pairs of vertices, alongside 2D visualizations. Each row is a different graph, and each column is a different learning rate. For both the learning rate $0.001$ and $0.01$, the geometries take about $200$ steps to appear; the larger learning rate of $0.1$ is too aggressive to create the geometry. Recall that, for associative memory, two steps with learning rate of $0.1$ sufficed (Fig.˜7). This is evidence for Section˜3.2: ease of discovery does not determine what sort of memory is preferred.

### v3

@@ B.2.2 Tiny model architectures
For toy experiments on small-scale graphs (§C.3), including the Tiny Path–Star, Tiny Grid, Tiny Cycle, and Tiny Irregular graphs, we used reduced-size architectures.
Models. We evaluated three model types: (1) a Transformer (TinyGPT), (2) a feed-forward linear neural network with the same configuration but without residual connection or attention (TinyNN) and (3) a Mamba model (TinySSM) of comparable scale to (1). Note that the TinyNN is initialized the same way as the TinyGPT: all trainable embedding and linear weights are initialized from $\mathcal{N}(0,0.02^{2})$, while linear biases are initialized to zero. With tied input/output embeddings, the language-model head shares the token embedding matrix and computes output logits via its transpose. Unless otherwise noted, all experiments use this GPT-style token embedding initialization. In some analyses (see §C.3.1), we also consider the PyTorch default token embedding initialization, where embeddings are sampled from $\mathcal{N}(0,1)$. Exact configurations of each of the models mentioned are described below.
@@
  • For all graphs, all models used a single layer ($N_{\tt layer}=1$) besides the embedding/unembedding layers, but the model specifications differed across graph types.
@@
  • For all tiny-graph experiments, TinyGPT used a single attention head ($m_{\tt head}=1$) and embedding dimension $m_{\tt width}=32$.
  • For Tiny Path–Star, Tiny Grid, and Tiny Cycle, TinySSM used embedding dimension $8$ with $d_{\text{state}}=2$, $d_{\text{conv}}=2$, and $\text{expand}=1$. For Tiny Irregular, TinySSM used embedding dimension $8$ with $d_{\text{state}}=2$, $d_{\text{conv}}=4$, and $\text{expand}=1$.
For TinyNN, we used a large embedding dimension $512$ with $1$ head for the sake of empirical proofs of Section˜3. We elaborate why. First, TinyNN architectures are a clean architecture where one can be certain that freezing the (un)embedding layers results in a purely associative storage; thus they are appropriate for our empirical proofs. Next, in our proofs, we wanted to make two demonstrations. First, that the model memorize geometrically even without bottleneck pressures (as is the case with a large embedding dimension); second, that in sufficiently wide-neck settings, the associative storage is realized within just $2$ learning steps (which we find is true here)
We clarify that even when reducing the embedding dimension of TinyNN to match that of TinyGPT (Tiny Path–Star and Tiny Grid: $m_{\tt width}=32$; Tiny Irregular: $m_{\tt width}=16$), we found that the geometric structure still persisted and that the associative memory variant continued to fully memorize the edge bigrams, albeit with longer optimization. In particular, Tiny Path–Star and Tiny Grid required roughly $500$ optimization steps with learning rate $0.01$, while Tiny Irregular required roughly $1000$ optimization steps with learning rate $0.005$. For Tiny Cycle, we continued to use a wider embedding dimension of $512$, which required roughly 10,00010{,}000 optimization steps at learning rate $0.01$. In contrast, in sufficiently wide settings (e.g., embedding dimension $512$), the associative storage mechanism is realized within just $2$ learning steps, which we find to hold both in these toy settings and in the associative path-star experiments of Section˜3..
@@
Two-layer (Node2Vec-style) architectures. For the shallow Node2Vec-style archirtectures (two weight-tied layers) used in §4, we use wide models with embedding dimension $m=100$ to ensure that the models have no explicit bottleneck/rank pressure. We train these models with a learning rate of $0.01$, a batch size equal to the number of nodes in the graph, and $2000$ steps, which is well beyond fitting the training data to full accuracy.
@@ B.3 Training and optimization
Objective. We use next-token cross-entropy over the causal prefix. For first-token-only experiments, the loss is restricted to the first target position. Although for large graph experiments we train all our models to 50,00050{,}000 epochs, we only need about a couple of thousand epochs to see accuracy gains (e.g., see the per-token accuracy plots in Fig.˜12).
@@
Tiny-graph optimization details. For the tiny-graph experiments in §C.3, optimization hyperparameters varied by graph type. For Tiny Path–Star and Tiny Grid, both TinyGPT and TinyNN were trained for $500$ optimization steps with learning rate $0.01$. For Tiny Cycle, TinyGPT was trained for $700$ steps with learning rate $0.01$, while TinyNN used embedding dimension $512$ and required 10,00010{,}000 optimization steps at learning rate $0.01$. For Tiny Irregular, TinyGPT was trained for $2000$ steps with learning rate $0.01$, while TinyNN used embedding dimension $16$, learning rate $0.005$, and $1000$ optimization steps.
For all experiments in Section˜C.3.1, we use a fixed training budget of 10,00010{,}000 optimization steps across all settings to ensure comparability between regularization strategies.
@@
B.4 Path-finding evaluation
@@
  • Per-token accuracy over epochs / Token accuracy: accuracy at each target position, tracked through training.
@@
B.5 Tests for (global) geometric vs associative storage
Path-finding is one way to test for the presence of a global geometry. But in our smaller models, we have used various simpler tests. These tests detect whether the embedding matrix indeed has information about the graph or whether it is arbitrary (if it is, then we know that the storage is associative at least in 3-layered TinyNN models).
Remark 4.
(Global vs. local geometry) If there is a geometry in the embeddings, we can ask whether it is global or relatively local e.g., is the embedding only close to its immediate neighbors but orthogonal to the rest? [^13] We wish to thank Omar Salemohamed for suggesting the possibility of such local geometries.). As a concrete example, consider the vertices of a path graph embedded along the edges of a hypercube as $(1,1,0,0,\ldots),(0,1,1,0,\ldots),(0,0,1,1,\ldots)$ and so on. Such a solution encodes a local geometry in that each vertex has high cosine-similarity to its neighbors, but to no other vertices. On this regard, each of our tests comes with its own false positive or negative as discussed below. We use the sum total of all tests to ascertain the presence of global geometries as against local geometries.
  1. 1.
    3D visualizations: We found 3D visualizations like in Fig.˜1 to help track many qualitative behaviors e.g., whether the embeddings are arbitrary, the presence of negative eigendirections in the form of zigzags, the varying quality of geometries under various interventions etc., However, we warn the reader that these visualizations can suffer from a false positive on the local/global test: models with local geometries could still display a global geometry in the 3D visualization since the visualization method may inject some global information into it.
  2. 2.
    Cosine heatmaps: In plots such as Figs.˜4 and 23, we provide heatmaps of the cosine distances between embeddings. If the heatmap resembles multi-hop adjacency, it affirms a global geometry; if it appears random, the embedding matrix has no useful information. However this can yield false negatives: if the heatmap resembles a spiky one-hop adjacency matrix, the geometry may still be global (which is possible if the embeddings are a full-rank factorization of the adjacency matrix).
  3. 3.
    Eigenvector projections: We also provide projections of the embeddings onto the eigenvectors of the graph, like in Fig.˜7. A random embedding matrix would be roughly equally distributed along all directions, whereas a more geometric embedding would reflect the underlying graph spectrum. In fact, a global geometry would correspond to a spectrum that is even more skewed towards the topmost directions (e.g., uses only the top two directions in the cycle). Our plots are designed as follows. We index the eigenvectors for the negative graph Laplacian by their signed eigenvalues. We ignore the top eigenvector which is a degenerate eigenvector that assigns near-uniform values to all nodes. Then, for various indices $i$, we plot the magnitude of $\|{\bf{V}}(t)^{T}{\bf{e}}_{i}\|/\|{\bf{V}}(t)^{T}{\bf{e}}_{1}\|$, which is the projection of the embeddings along the $i$’th direction relative to the projection along the second-top direction. We compare the spread of these relative magnitudes against that of the eigenvalues themselves as $\lambda_{i}/\lambda_{1}$.
  4. 4.
    Geometric margin: As a more quantitative test for geometries, for our 3-layered TinyNN model, we compute the fit of a modified model that is purely geometric. We do this by eliminating all cross-eigenvector connections in the middle layer $\bf{W}$}. Specifically, if $\hat{{\bf{V}}}$ are the right-singular directions of the learned embedding matrix $\hat{{\bf{\Phi}}}$, we replace the middle layer with $\hat{{\bf{V}}}^{T}\texttt{diag}(\hat{{\bf{V}}}{\bf{W}}\hat{{\bf{V}}}^{T})\hat{{\bf{V}}}$. We then compute the margin for this modified model on the training set as $\min_{u}\{\min_{w\in\texttt{nbr}(u)}f(u)[w]-\max_{v\notin\texttt{nbr}(u)}f(u)[v]\}$. If this is positive, it means that the embedding matrices encode the graph structure through cosine-similarity alone (modulo some sign and magnitudes prescribed by the diagonal values in the intermediate layer). If the geometric margin is negative however, it may be a false negative because there may be different diagonal values that result in a fit of the data.
B.6 Implementation and compute
Code is implemented in PyTorch with standard Transformer components. All runs fit on a single modern GPU (e.g., A100-40GB or A100-80GB for the largest batch-size); per-figure training time depends on $(d,\ell)$ and batch size. Code for reproducing the experiments is available at: https://github.com/shahriarnz14/geometric_memory.
@@ C.1 Path-star task on Mamba SSM
This section (Figs.˜14 and 15) demonstrates that the implicit reasoning on path-star graphs observed for Transformers in §2.1, generalizes to the Mamba SSM architecture [Gu and Dao, 2023] too. A notable difference here is that the Mamba model presents a strong geometry even when trained only on the edges (Fig.˜15(b)).
Figure 14: (left) Success of in-weights path-star task for Mamba. This figure is a counterpart to Fig.˜12. A next-token-trained Mamba achieves perfect or highly non-trivial accuracy on large path-star graphs $\mathcal{G}_{d,\ell}$ (Section˜A.1). (middle) Learning order of tokens. The tokens of a path are not learned in the reverse order i.e., the model does not learn the right-to-left solution. Thus, gradients from the future tokens are not critical for success (Section˜A.1.1). (right) Success of hardest-token-only task. In fact, the hardest token (the first token) given the leaf is learned in isolation to non-trivial accuracy (Section˜2.2). Success of this $\ell$-fold composition task is hard to explain within the associative memory (§2.3).
@@
(b) Only Edge-Memorized
Figure 15: Evidence of global geometry in path-star task for Mamba. This figure is a Mamba counterpart to the Transformer heatmaps in Figs.˜5 and 4. Recall that entry $(i,j)$ is the mean cosine distance between the leaf token in (an unseen) path $i$ (row) and first/hardest token on (an unseen) path $j$ (col). Left is trained on edges and path-finding task $(\mathcal{D}_{\texttt{edge}}\cup\mathcal{D}_{\texttt{path}}^{\to})$; right on edges only $(\mathcal{D}_{\texttt{edge}})$. We find that even on these unseen paths, the leaf and first token embeddings cluster together, regardless of whether path-finding supervision exists. Interestingly, compared to the Transformer in Fig.˜5, the Mamba trained only on local supervision (right) exhibits a much stronger geometry.
@@ C.2 Other large, harder path-finding graphs
Next, we consider variants of the path-finding graph. While the path-star graph is adversarially constructed in certain ways, there is only one decision-making step, which makes planning simpler in a certain way. To make the task harder, we introduce a branching at every node along each path. (Similar tree variants have been considered in [Frydenlund, 2025] for the in-context task, but we are interested in in-weights tasks.)
In the tree-star graph, $\mathcal{T}_{d,\ell}$, there is a central node with degree $d$ and each child node except the leaf has $2$ children, as visualized in Fig.˜16(a). The path-length from the root to any leaf is $\ell$. In this graph, two types of learning tasks can be considered, depending on how we split the test and training paths. For a no-overlap setting, we could sample all training paths from one subset of trees, and the test paths from the remaining trees; we call this the split at first token setting, since the test/train split is determined by the first token. A second setting—with some test-train overlap—is one where we reserve some leaves as training goals, and the rest as test-goals. Here, on any test path, a prefix may have participated in a training path. We call this the split at leaf setting. Note that in both variants, all nodes will be sampled as part of the edge-memorization task.
In Fig.˜16(b), we find that on both variants the Transformer achieves non-trivial path-finding accuracy, generalizing our results beyond the path-star task. However, we note that the test-train split at the first token is much harder to succeed at, likely due to no overlaps.
@@ C.3 Tiny graphs
Graphs Besides the large path-star graph (of §2) and tree-star graphs (of C.2), we also report the embeddings learned on four tinier graphs for various architectures (details in B.2.2); the figures here are an extension of the Transformer embeddings in Fig.˜1. In these experiments, we train the models purely on local supervision (the edges, presented in both directions) to $100\%$ edge-memorization accuracy: for each vertex, we ensure that its $d$ neighbors appear in the top $d$ softmax probabilities. The graphs include (a) a tiny path-star graph with four paths of length $4$, (b) a $4\times 4$ grid graph, (c) a $15$-node cycle graph and (d) an irregular graph with two asymmetric components.
List of figures. Our visualization techniques are based on the tests described in §B.5. Our first set of figures (Figs.˜17, 18, 19 and 20) are 3D visualizations of the embeddings from various architectures: an associative memory model (implemented with a neural network with only one trainable matrix sandwiched between (un)embedding layers), a Node2Vec model, the eigenvectors of the graph Laplacian, a Transformer’s token embeddings, a neural network’s first layer, and a Mamba SSM’s token embeddings. For the Node2Vec model we use the top eigenvectors, and for the rest we use UMAP [McInnes et al., 2018] to choose the top directions. As noted in §B.5, these visualizations by themselves do not guarantee the presence of global information. As a rigorous test, in Figs.˜24 and 23, we provide node-node cosine-similarity heatmaps exhibiting multi-hop information. Next in Figs.˜22 and 21, we present detailed plots of how fast a model stores associatively/geometrically over training time. In §C.3.1, we provide plots demonstrating the effect of various optimization hyperparameters.
Consolidated Observations. For easy reference, we consolidate all our observations from these plots below:
In the tiny graphs of Figs.˜17, 18, 19, 20, 22, 21, 24 and 23 and §C.3.1 on various architectures (Node2Vec, graph spectrum, Transformer, neural network, Mamba SSM), we find that:
@@
    While the geometry is significantly aided by higher weight decay or dropout, these regularizing effects are not necessary (Section˜3.2); (demonstrated in §C.3.1).
  5. 5.
@@
    Associative storage can be quickly found by gradient descent (Fig.˜21), as claimed in Section˜3.2, whereas geometric memorization takes longer to discover (Fig.˜22)
  7. 7.
    Geometries arise even with only one direction of the edges presented (Fig.˜32).
@@
Figure 20: Tiny irregular graph: Geometries of various architectures on a small irregular graph of two connected components, both asymmetric. See Section˜C.3.
(a) Tiny Path-Star
(b) Tiny Grid
(c) Tiny Cycle
(d) Tiny Irregular
Figure 21: Associative memory can be discovered quickly by gradient descent, given a sufficiently wide model, and a sufficiently large learning rate: For the various tiny graphs described in §C.3, and for our TinyNN model (with frozen embedding/unembedding layers and one wide trainable weight matrix to prevent a geometry from taking over; see §B.2.2), we report memorization over timesteps of training with full-batch gradient descent and constant learning rate of $0.1$. For each vertex, we compute what fraction of its $k$ neighbors appear in the top $k$ next token probabilities; this is averaged across vertices (where $k$ varies across vertices). This value, which quantifies associative memorization, reaches its maximum within 22 steps of gradient descent; compare this to the longer time for geometric memory to form in Fig.˜22. This is evidence for our Section˜3.2: ease-of-discovery does not dictate what memory is formed.
@@
Figure 22: For various learning rates, geometric memorization takes much longer for gradient descent to discover: This is an extended version of Fig.˜6(b) for two more learning rates and all our tiny graphs. Recall that for the same TinyNN architecture as in Fig.˜21 but without frozen embeddings, we roughly quantify geometric memorization over time via cosine distance across all pairs of vertices, alongside 2D visualizations. Each row is a different graph, and each column is a different learning rate. For both the learning rate $0.001$ and $0.01$, the geometries take about $200$ steps to appear; the larger learning rate of $0.1$ is too aggressive to create the geometry. Recall that, for associative memory, two steps with learning rate of $0.1$ sufficed (Fig.˜21). This is evidence for Section˜3.2: ease of discovery does not determine what sort of memory is preferred.